OpenStudy (anonymous):

Find a formula for the inverse of the function. f(x) = 6x − 1/2x + 7

OpenStudy (amistre64):

is this typed correctly?

OpenStudy (anonymous):

6x-1 divide by 2x+7 sorry

OpenStudy (amistre64):

:)

OpenStudy (amistre64):

change x to y and solve for y tends to work

OpenStudy (anonymous):

i got stuck once i multiplied x(2y+7)

OpenStudy (anonymous):

2xy +7x+1=6y now what

OpenStudy (amistre64):

6y -1 ----- = x 2y+7 6y -1 = x(2y+7) 6y = 2xy +7x +1 6y - 2xy = 7x +1 y(6-2x) = 7x +1 7x+1 y = ----- 6-2x

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

yw :)